OpenStudy (anonymous):

A coin is tossed six times. What is the probability of obtaining at most five heads?

OpenStudy (anonymous):

1/64

OpenStudy (anonymous):

For every toss, you have "1/2" probability of getting heads. So for 6 tosses, you have (1/2)^6 or 1/64

OpenStudy (anonymous):

P(at most 5 heads) = 1 - P(6 heads) = 1 - (1/2)^6 = 63/64

OpenStudy (anonymous):

the questions is asking for AT MOST 5 heads, NOT exactly 5 heads

OpenStudy (anonymous):

Overlooked the problem... sorry for the confusion

OpenStudy (anonymous):

Thank you